OpenStudy (anonymous):

Find the derivative of f(x) = negative 6 divided by x at x = 12.

OpenStudy (anonymous):

\[f(x) = - \frac{ 6 }{ x }\] this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Alright so any idea how to find the derivative? Hint: Power rule

OpenStudy (anonymous):

No I am new to this lesson

OpenStudy (anonymous):

Do you know of the power rule?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

\[\frac{ d }{ dx } x^n \implies n x^{n-1}\]

OpenStudy (anonymous):

Power rule, as you can see you just bring the exponent down, and subtract by 1 in the exponents

OpenStudy (anonymous):

So for an example, if I say find the derivative of \[x^9\] all you'd do is \[f'(x) = 9x^8\]

OpenStudy (anonymous):

so would it be f(x)=2x^(-6/x-1)??

OpenStudy (anonymous):

Mhm, not sure what you did there, where did 2x come from?

OpenStudy (anonymous):

Sorry I meant 12 because it says at x=12

OpenStudy (anonymous):

Well what I suggest is, you find the derivative first then plug in 12 for x.

OpenStudy (anonymous):

Ohh ok so f(x)=x^-6x-1?

OpenStudy (anonymous):

No, that's not right. The calculus part is easy, the main thing here is knowing your exponent rules. What can we do with \[\frac{ 1 }{ x }\]

OpenStudy (anonymous):

If I asked you to do use the power rule, how would you do it, just with 1/x

OpenStudy (anonymous):

Think back to exponents

OpenStudy (anonymous):

Ok I would do f(x)x^(1/x-1)

OpenStudy (anonymous):

oh wait is it 1/x^(1/x-1)?

OpenStudy (anonymous):

No no, you're over thinking this, note that \[\frac{ 1 }{ x } \implies x^{-1}\] right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So the derivative would just be, \[\large -x^{-2} \implies -\frac{ 1 }{ x^2 }\]

OpenStudy (anonymous):

Ohhh ok that makes sense

OpenStudy (anonymous):

Yup! Now what would be the derivative of \[f(x) = -\frac{ 6 }{ x }\]

OpenStudy (anonymous):

f(x)=-6x^-7

OpenStudy (anonymous):

-7?

OpenStudy (anonymous):

I think you need to review your exponent rules, which I made a very short tutorial for, you can find it here: http://openstudy.com/users/iambatman#/updates/54ff9b2ce4b064b5a94d524c

OpenStudy (anonymous):

Ok Im currently looking. I might take a while but I'll message you back

OpenStudy (anonymous):

Np, take as much time as you like, I don't want to give you the answer because you will not learn that way, so when ever you're ready go ahead and post what you think it is :) Good luck.

OpenStudy (anonymous):

Ok I'm back so the derivative would be 6/x^2

OpenStudy (anonymous):

Good!

OpenStudy (anonymous):

Finally! what do I do with the x at 12 part?

OpenStudy (anonymous):

Now you have the derivative, you simply plug it in for x :)

OpenStudy (anonymous):

Oh! so the answer is 1/24!

OpenStudy (anonymous):

Yup :)!

OpenStudy (anonymous):

Yay! Thank you so much @iambatman for helping me :)

OpenStudy (anonymous):

No problem, take care!